OpenStudy (anonymous):

solve 1 + 1 1 -- -- = -- x+3 x-3 x^-9

jhonyy9 (jhonyy9):

1/x+3 +1/x-3 =1/x*2 -9 x-3 + x+3 = 1 2x=1 x=1/2